# Matrix Manifold Neural Networks++

### v1

@@ D.4.1 Ablation Study
To study the impact of the factor $\beta$ in Affine-Invariant metrics on the performance of GyroSpd++, we follow the approach in Nguyen (2021). Denote by $(\boldsymbol{\mu},\boldsymbol{\Sigma})$ the Gaussian distribution where $\boldsymbol{\mu}\in\mathbb{R}^{n}$ and $\boldsymbol{\Sigma}\in\operatorname{M}_{n,n}$ are its mean and covariance. We can identify the Gaussian distribution $(\boldsymbol{\mu},\boldsymbol{\Sigma})$ with the following matrix:
@@
where $k\geq 1$, $\boldsymbol{\mu}(k)$ is a matrix with $k$ identical column vectors $\boldsymbol{\mu}$. The natural symmetric Riemannian metric resulting from the above embedding is given (Nguyen, 2021) by
@@ D.4.2 Comparison of GyroSpd++ against State-of-the-Art Methods
Here we present more comparisons of our networks against state-of-the-art networks. These networks belong to one of the following families of neural networks: (1) Hyperbolic neural networks: HypGRU (Ganea et al., 2018) [^3] https://github.com/dalab/hyperbolic_nn.; (2) Graph neural networks: MS-G3D (Liu et al., 2020) [^4] https://github.com/kenziyuliu/MS-G3D., TGN (Zhou et al., 2023) [^5] https://github.com/zhysora/FR-Head.; (3) Transformers: ST-TR (Plizzari et al., 2021) [^6] https://github.com/Chiaraplizz/ST-TR.. Note that MS-G3D, TGN, and ST-TR are specifically designed for skeleton-based action recognition. We use default parameter settings for these networks. Results of our networks and their competitors on HDM05, FPHA, and NTU60 datasets are shown in Tabs. 78, and 9, respectively. On HDM05 dataset, GyroSpd++ outperforms HypGRU, MS-G3D, ST-TR, and TGN by 25.6%, 10.8%, 11.9%, and 9.5% points in terms of mean accuracy, respectively. On FPHA dataset, GyroSpd++ outperforms HypGRU, MS-G3D, ST-TR, and TGN by 38.6%, 8.5%, 10.9%, and 6.0% points in terms of mean accuracy, respectively. On NTU60 dataset, GyroSpd++ outperforms HypGRU, MS-G3D, ST-TR, and TGN by 7.0%, 3.1%, 4.2%, and 2.2% points in terms of mean accuracy, respectively. Overall, our networks are superior to their competitors in all cases.
@@
HypGRU (Ganea et al., 2018) 54.18 $\pm$ 1.51
MS-G3D (Liu et al., 2020) 68.92 $\pm$ 1.72
ST-TR (Plizzari et al., 2021) 67.84 $\pm$ 1.66
TGN (Zhou et al., 2023) 70.26 $\pm$ 1.48
@@
HypGRU (Ganea et al., 2018) 58.24 $\pm$ 0.29
MS-G3D (Liu et al., 2020) 88.26 $\pm$ 0.67
ST-TR (Plizzari et al., 2021) 85.94 $\pm$ 0.46
TGN (Zhou et al., 2023) 90.81 $\pm$ 0.53
@@
HypGRU (Ganea et al., 2018) 88.26 $\pm$ 1.40
MS-G3D (Liu et al., 2020) 92.15 $\pm$ 1.60
ST-TR (Plizzari et al., 2021) 91.04 $\pm$ 1.52
TGN (Zhou et al., 2023) 93.02 $\pm$ 1.56
@@ E.2.1 Setup
Our network is implemented using the PyTorch framework. We set hyperparameters as in Zhao et al. (2023) that are found via grid search for each graph architecture on the development set of a given dataset. The best settings of $n$ and $p$ are found from $(n,p)\in\{(2k,k)\},k=2,3,\ldots,10$. The batch size is set to the total number of graph nodes in a dataset (Chami et al., 2019; Zhao et al., 2023). The networks are trained using cross-entropy loss and Adam optimizer for a maximum of 500 epochs. The learning rate is set to $10^{-2}$. Early stopping is used when the loss on the development set has not decreased for 200 epochs. Each network has two layers that perform message passing twice at one iteration (Zhao et al., 2023). We use the 70/15/15 percent splits (Chami et al., 2019) for Airport dataset, and standard splits in GCN Kipf & Welling (2017) for Pubmed and Cora datasets.
@@ E.2.2 Grassmann logarithmic map in the ONB perspective
The Grassmann logarithmic map in the ONB perspective is given (Edelman et al., 1998) by
@@ E.2.3 Gr-GCN++
To create Grassmann embeddings as input node features, we first transform $d$-dimensional input features into $p(n-p)$-dimensional vectors via a linear map. We then reshape each resulting vector to a matrix $\mathbf{B}\in\operatorname{M}_{p,n-p}$. The input Grassmann embedding $\mathbf{X}_{i}^{0},i\in\mathcal{V}$ is computed as
$\mathbf{X}_{i}^{0}=\exp\bigg{(}\begin{bmatrix}0&\mathbf{B}\\ -\mathbf{B}^{T}&0\end{bmatrix}\bigg{)}\mathbf{I}_{n,p}\exp\bigg{(}-\begin{bmatrix}0&\mathbf{B}\\ -\mathbf{B}^{T}&0\end{bmatrix}\bigg{)}.$
@@ E.2.4 Gr-GCN-ONB
To create Grassmann embeddings as input node features, we first transform $d$-dimensional input features into $p(n-p)$-dimensional vectors via a linear map. We then reshape each resulting vector to a matrix $\mathbf{B}\in\operatorname{M}_{p,n-p}$. The input Grassmann embedding $\mathbf{X}_{i}^{0},i\in\mathcal{V}$ is computed as
@@
Feature transformation is performed by first mapping the input to a projection matrix (using the mapping $\tau$ in Section 3.4.1), then applying an isometry map based on left Grassmann gyrotranslations (Nguyen & Yang, 2023), and finally mapping the result back to a matrix with orthonormal columns. This is equivalent to performing the following mapping:
@@
Bias addition is performed using operation ~$\widetilde{\oplus}_{gr}$ instead of operation $\oplus_{gr}$. The output of Gr-GCN-ONB is mapped to a projection matrix for node classification.
@@ E.2.5 Optimization
$\mathbf{P}=\exp\bigg{(}\begin{bmatrix}0&\mathbf{B}\\ -\mathbf{B}^{T}&0\end{bmatrix}\bigg{)}\mathbf{I}_{n,p}\exp\bigg{(}-\begin{bmatrix}0&\mathbf{B}\\ -\mathbf{B}^{T}&0\end{bmatrix}\bigg{)}.$
@@ E.3.1 Ablation Study
GCN (Kipf & Welling, 2017) 82.2 $\pm$ 0.6 77.8 $\pm$ 0.8 80.2 $\pm$ 2.3
GAT (Velic̆ković et al., 2018) 92.9 $\pm$ 0.8 77.6 $\pm$ 0.8 80.3 $\pm$ 0.6
HGNN (Liu et al., 2019) 84.5 $\pm$ 0.7 76.6 $\pm$ 1.4 79.5 $\pm$ 0.9
HGCN (Chami et al., 2019) 85.3 $\pm$ 0.6 76.4 $\pm$ 0.8 78.7 $\pm$ 0.9
LGCN (Zhang et al., 2021) 88.2 $\pm$ 0.2 77.3 $\pm$ 1.4 80.6 $\pm$ 0.9
HGAT (Zhang et al., 2022) 87.5 $\pm$ 0.9 78.0 $\pm$ 0.5 80.9 $\pm$ 0.7
SPD-GCN (Zhao et al., 2023) 82.6 $\pm$ 1.5 78.7 $\pm$ 0.5 82.3 $\pm$ 0.5
HamGNN (Kang et al., 2023) 95.9 $\pm$ 0.1 78.3 $\pm$ 0.6 80.1 $\pm$ 1.6
@@ E.3.2 Comparison of Gr-GCN++ against State-of-the-Art Methods
Tab. 14 shows results of Gr-GCN++ and some state-of-the-art methods on the three datasets. The hyperbolic networks outperform their SPD and Grassmann counterparts on Airport dataset with high hyperbolicity (Chami et al., 2019). This agrees with previous works (Chami et al., 2019; Zhang et al., 2022) that report good performances of hyperbolic embeddings on tree-like datasets. However, our network and its SPD counterpart SPD-GCN outperform their competitors on Pubmed and Cora datasets with low hyperbolicities. Compared to SPD-GCN, Gr-GCN++ always gives more consistent results.
@@ Appendix F Limitations of Our Work
In our Grassmann network Gr-GCN++, the feature transformation and bias and nonlinearity operations are performed on Grassmann manifolds, while the aggregation operation is performed in tangent spaces. Previous works (Dai et al., 2021; Chen et al., 2022) on HNNs have shown that this hybrid method limits the modeling ability of networks. Therefore, it is desirable to develop GCNs where all the operations are formalized on Grassmann manifolds.
@@ G.1 Gyrogroups and Gyrovector Spaces
Gyrovector spaces form the setting for hyperbolic geometry in the same way that vector spaces form the setting for Euclidean geometry (Ungar, 2002; 2005; 2014). We recap the definitions of gyrogroups and gyrocommutative gyrogroups proposed in Ungar (2002; 2005; 2014). For greater mathematical detail and in-depth discussion, we refer the interested reader to these papers.
Definition G.1 (Gyrogroups (Ungar, 2014)).
@@
[…]
@@
Definition G.2 (Gyrocommutative Gyrogroups (Ungar, 2014)).
@@
[…]
The following definition of gyrovector spaces is slightly different from Definition 3.2 in Ungar (2014).
@@ G.2 AI Gyrovector Spaces
For $\mathbf{P},\mathbf{Q}\in\operatorname{Sym}_{n}^{+}$, the binary operation (Nguyen, 2022a) is given as
@@
The inverse operation (Nguyen, 2022a) is given by
@@ G.3 LE Gyrovector Spaces
For $\mathbf{P},\mathbf{Q}\in\operatorname{Sym}_{n}^{+}$, the binary operation (Nguyen, 2022a) is given as
@@
The inverse operation (Nguyen, 2022a) is given as
@@ G.4 LC Gyrovector Spaces
For $\mathbf{P},\mathbf{Q}\in\operatorname{Sym}_{n}^{+}$, the binary operation (Nguyen, 2022a) is given as
@@
The inverse operation (Nguyen, 2022a) is given by
@@ G.5 Grassmann Manifolds in The Projector Perspective
For $\mathbf{P},\mathbf{Q}\in\operatorname{Gr}_{n,p}$, the binary operation (Nguyen, 2022b) is given as
@@
where [.,.][.,.][ . , . ] denotes the matrix commutator.
The inverse operation (Nguyen, 2022b) is defined as
@@ G.6 Grassmann Manifolds in The ONB Perspective
For $\mathbf{U},\mathbf{V}\in\widetilde{\operatorname{Gr}}_{n,p}$, the binary operation (Nguyen & Yang, 2023) is defined as
@@
The inverse operation can be defined using the approach in Nguyen & Yang (2023) (see Section 2.3.1), i.e.,
~$\widetilde{\ominus}_{gr}\mathbf{U}=\tau^{-1}\big{(}\ominus_{gr}(\mathbf{U}\mathbf{U}^{T})\big{)},$
@@ G.7 The SPD and Grassmann Inner Products
where .,.$\langle.,.\rangle_{\mathbf{I}_{n,p}}$ denotes the inner product at $\mathbf{I}_{n,p}$ given by the canonical metric of $\operatorname{Gr}_{n,p}$.
@@ G.8 The Gyrocosine Function and Gyroangles in Structure Spaces
where .$\|.\|^{psd,g}$ is the norm induced by the inner product in structure spaces. The gyroangle $\alpha$ is denoted by $\alpha=\angle\mathbf{Q}\mathbf{P}\mathbf{R}$.
@@ Appendix H Computation of Canonical Representations
We first recall the definition of the binary operation $\oplus_{g}$ in Nguyen (2022b).
Definition I.1 (The Binary Operation (Nguyen, 2022b)).
@@
The first part of Proposition 3.4 can be easily verified using the definition of the SPD inner product (see Definition G.4) and that of Affine-Invariant metrics (Pennec et al., 2020) (see Chapter 3).
To prove the second part of Proposition 3.4, we will use the notion of SPD pseudo-gyrodistance (Nguyen & Yang, 2023) in our interpretation of FC layers on SPD manifolds, i.e., the signed distance is replaced with the signed SPD pseudo-gyrodistance in the interpretation given in Section 3.2.1. First, we need the following result from Nguyen & Yang (2023).
Theorem J.1 (The SPD Pseudo-gyrodistance from an SPD Matrix to an SPD Hypergyroplane in an AI Gyrovector Space (Nguyen & Yang, 2023)).
@@
[…]
@@
[…]
@@
The first part of Proposition 3.6 can be easily verified using the definition of the SPD inner product (see Definition G.4) and that of Log-Cholesky metrics (Lin, 2019).
To prove the second part of Proposition 3.6, we first recall the following result from Nguyen & Yang (2023).
Theorem L.1 (The SPD Gyrodistance from an SPD Matrix to an SPD Hypergyroplane in a LC Gyrovector Space (Nguyen & Yang, 2023)).
@@
$\mathbf{B}=\lfloor\widetilde{\mathbf{W}}\rfloor+\mathbb{D}(\varphi(\mathbf{P}))^{-1}\mathbb{D}(\widetilde{\mathbf{W}}),$
@@
[…]
@@
Now, the problem in (9) is to find the minimum angle between the vector $[\sqrt{\lambda}\mathbf{A}_{1}\|\mathbf{A}_{2}]$ and the Euclidean hyperplane described by Eq. (11). The pseudo-gyrodistance from $\mathbf{X}$ to $\mathcal{H}_{\mathbf{W},\mathbf{P}}^{psd,g}$ thus can be obtained as
@@
We need the following result from Nguyen & Yang (2023).
@@
where $\mathbf{P},\mathbf{Q}\in M$, $D\tau_{\mathbf{R}}(\mathbf{W})$ denotes the directional derivative of a mapping $\tau$ at point $\mathbf{R}\in N$ along direction $\mathbf{W}\in T_{\mathbf{R}}N$, $\operatorname{Log}(.)$ and $\widetilde{\operatorname{Log}}(.)$ are the logarithmic maps in manifolds M and N, respectively.
We adopt the notations in Bendokat et al. (2020). The Riemannian metric $g^{O}_{\mathbf{Q}}(.,.)$ on $\operatorname{O}_{n}$ is the standard inner product given (Edelman et al., 1998; Bendokat et al., 2020) as
@@
Let $\mathbf{U}\in\widetilde{\operatorname{Gr}}_{n,p}$, $\mathbf{D}_{1},\mathbf{D}_{2}\in T_{\mathbf{U}}\widetilde{\operatorname{Gr}}_{n,p}$. The canonical metric $g^{\widetilde{\operatorname{Gr}}}_{\mathbf{U}}(\mathbf{D}_{1},\mathbf{D}_{2})$ on $\widetilde{\operatorname{Gr}}_{n,p}$ is the restriction of the Riemannian metric $g^{O}_{\mathbf{Q}}(.,.)$ to the horizontal space of $T_{\mathbf{Q}}\operatorname{O}_{n}$ (multiplied by 1/2) and is given (Edelman et al., 1998; Bendokat et al., 2020) by
@@
Let $\mathbf{P}\in\operatorname{Gr}_{n,p}$, $\Delta_{1},\Delta_{2}\in T_{\mathbf{P}}\operatorname{Gr}_{n,p}$. The canonical metric $g^{Gr}_{\mathbf{P}}(\Delta_{1},\Delta_{2})$ on $\operatorname{Gr}_{n,p}$ is the restriction of the Riemannian metric $g^{O}_{\mathbf{Q}}(.,.)$ to the horizontal space of $T_{\mathbf{Q}}\operatorname{O}_{n}$ (multiplied by 1/2) and is given (Edelman et al., 1998; Bendokat et al., 2020) by

### v2

@@ D.4.1 Ablation Study
To study the impact of the factor $\beta$ in Affine-Invariant metrics on the performance of GyroSpd++, we follow the approach in Nguyen_2021_ICCV. Denote by $(\boldsymbol{\mu},\boldsymbol{\Sigma})$ the Gaussian distribution where $\boldsymbol{\mu}\in\mathbb{R}^{n}$ and $\boldsymbol{\Sigma}\in\operatorname{M}_{n,n}$ are its mean and covariance. We can identify the Gaussian distribution $(\boldsymbol{\mu},\boldsymbol{\Sigma})$ with the following matrix:
@@
where $k\geq 1$, $\boldsymbol{\mu}(k)$ is a matrix with $k$ identical column vectors $\boldsymbol{\mu}$. The natural symmetric Riemannian metric resulting from the above embedding is given (Nguyen_2021_ICCV) by
@@ D.4.2 Comparison of GyroSpd++ against State-of-the-Art Methods
Here we present more comparisons of our networks against state-of-the-art networks. These networks belong to one of the following families of neural networks: (1) Hyperbolic neural networks: HypGRU (NEURIPS2018_dbab2adc) [^3] https://github.com/dalab/hyperbolic_nn.; (2) Graph neural networks: MS-G3D (liu2020disentangling) [^4] https://github.com/kenziyuliu/MS-G3D., TGN (zhou2023learning) [^5] https://github.com/zhysora/FR-Head.; (3) Transformers: ST-TR (Plizzari2021skeleton) [^6] https://github.com/Chiaraplizz/ST-TR.. Note that MS-G3D, TGN, and ST-TR are specifically designed for skeleton-based action recognition. We use default parameter settings for these networks. Results of our networks and their competitors on HDM05, FPHA, and NTU60 datasets are shown in Tabs. 78, and 9, respectively. On HDM05 dataset, GyroSpd++ outperforms HypGRU, MS-G3D, ST-TR, and TGN by 25.6%, 10.8%, 11.9%, and 9.5% points in terms of mean accuracy, respectively. On FPHA dataset, GyroSpd++ outperforms HypGRU, MS-G3D, ST-TR, and TGN by 38.6%, 8.5%, 10.9%, and 6.0% points in terms of mean accuracy, respectively. On NTU60 dataset, GyroSpd++ outperforms HypGRU, MS-G3D, ST-TR, and TGN by 7.0%, 3.1%, 4.2%, and 2.2% points in terms of mean accuracy, respectively. Overall, our networks are superior to their competitors in all cases.
@@
HypGRU (NEURIPS2018_dbab2adc) 54.18 $\pm$ 1.51
MS-G3D (liu2020disentangling) 68.92 $\pm$ 1.72
ST-TR (Plizzari2021skeleton) 67.84 $\pm$ 1.66
TGN (zhou2023learning) 70.26 $\pm$ 1.48
@@
HypGRU (NEURIPS2018_dbab2adc) 58.24 $\pm$ 0.29
MS-G3D (liu2020disentangling) 88.26 $\pm$ 0.67
ST-TR (Plizzari2021skeleton) 85.94 $\pm$ 0.46
TGN (zhou2023learning) 90.81 $\pm$ 0.53
@@
HypGRU (NEURIPS2018_dbab2adc) 88.26 $\pm$ 1.40
MS-G3D (liu2020disentangling) 92.15 $\pm$ 1.60
ST-TR (Plizzari2021skeleton) 91.04 $\pm$ 1.52
TGN (zhou2023learning) 93.02 $\pm$ 1.56
@@ E.2.1 Setup
Our network is implemented using the PyTorch framework. We set hyperparameters as in zhao2023modeling that are found via grid search for each graph architecture on the development set of a given dataset. The best settings of $n$ and $p$ are found from $(n,p)\in\{(2k,k)\},k=2,3,\ldots,10$. The batch size is set to the total number of graph nodes in a dataset (chami2019hyperbolic; zhao2023modeling). The networks are trained using cross-entropy loss and Adam optimizer for a maximum of 500 epochs. The learning rate is set to $10^{-2}$. Early stopping is used when the loss on the development set has not decreased for 200 epochs. Each network has two layers that perform message passing twice at one iteration (zhao2023modeling). We use the 70/15/15 percent splits (chami2019hyperbolic) for Airport dataset, and standard splits in GCN kipf2017semisupervised for Pubmed and Cora datasets.
@@ E.2.2 Grassmann logarithmic map in the ONB perspective
The Grassmann logarithmic map in the ONB perspective is given (Edelman98) by
@@ E.2.3 Gr-GCN++
To create Grassmann embeddings as input node features, we first transform $d$-dimensional input features into p(np)p(n-p)-dimensional vectors via a linear map. We then reshape each resulting vector to a matrix $\mathbf{B}\in\operatorname{M}_{p,n-p}$. The input Grassmann embedding $\mathbf{X}_{i}^{0},i\in\mathcal{V}$ is computed as
[…]
@@ E.2.4 Gr-GCN-ONB
To create Grassmann embeddings as input node features, we first transform $d$-dimensional input features into p(np)p(n-p)-dimensional vectors via a linear map. We then reshape each resulting vector to a matrix $\mathbf{B}\in\operatorname{M}_{p,n-p}$. The input Grassmann embedding $\mathbf{X}_{i}^{0},i\in\mathcal{V}$ is computed as
@@
Feature transformation is performed by first mapping the input to a projection matrix (using the mapping $\tau$ in Section 3.4.1), then applying an isometry map based on left Grassmann gyrotranslations (NguyenGyroMatMans23), and finally mapping the result back to a matrix with orthonormal columns. This is equivalent to performing the following mapping:
@@
Bias addition is performed using operation $\widetilde{\oplus}_{gr}$ instead of operation $\oplus_{gr}$. The output of Gr-GCN-ONB is mapped to a projection matrix for node classification.
@@ E.2.5 Optimization
[…]
@@ E.3.1 Ablation Study
GCN (kipf2017semisupervised) 82.2 $\pm$ 0.6 77.8 $\pm$ 0.8 80.2 $\pm$ 2.3
GAT (velickovic2018graph) 92.9 $\pm$ 0.8 77.6 $\pm$ 0.8 80.3 $\pm$ 0.6
HGNN (liu2019hgnn) 84.5 $\pm$ 0.7 76.6 $\pm$ 1.4 79.5 $\pm$ 0.9
HGCN (chami2019hyperbolic) 85.3 $\pm$ 0.6 76.4 $\pm$ 0.8 78.7 $\pm$ 0.9
LGCN (zhang2021lgcn) 88.2 $\pm$ 0.2 77.3 $\pm$ 1.4 80.6 $\pm$ 0.9
HGAT (zhang2022hgan) 87.5 $\pm$ 0.9 78.0 $\pm$ 0.5 80.9 $\pm$ 0.7
SPD-GCN (zhao2023modeling) 82.6 $\pm$ 1.5 78.7 $\pm$ 0.5 82.3 $\pm$ 0.5
HamGNN (kang2023hamilton) 95.9 $\pm$ 0.1 78.3 $\pm$ 0.6 80.1 $\pm$ 1.6
@@ E.3.2 Comparison of Gr-GCN++ against State-of-the-Art Methods
Tab. 14 shows results of Gr-GCN++ and some state-of-the-art methods on the three datasets. The hyperbolic networks outperform their SPD and Grassmann counterparts on Airport dataset with high hyperbolicity (chami2019hyperbolic). This agrees with previous works (chami2019hyperbolic; zhang2022hgan) that report good performances of hyperbolic embeddings on tree-like datasets. However, our network and its SPD counterpart SPD-GCN outperform their competitors on Pubmed and Cora datasets with low hyperbolicities. Compared to SPD-GCN, Gr-GCN++ always gives more consistent results.
@@ Appendix F Limitations of Our Work
In our Grassmann network Gr-GCN++, the feature transformation and bias and nonlinearity operations are performed on Grassmann manifolds, while the aggregation operation is performed in tangent spaces. Previous works (dai2021hhgcn; chen-etal-2022-fully) on HNNs have shown that this hybrid method limits the modeling ability of networks. Therefore, it is desirable to develop GCNs where all the operations are formalized on Grassmann manifolds.
@@ G.1 Gyrogroups and Gyrovector Spaces
Gyrovector spaces form the setting for hyperbolic geometry in the same way that vector spaces form the setting for Euclidean geometry (UngarBeyonEinstein02; UngarAnalyticHyp05; UngarHyperbolicNDim). We recap the definitions of gyrogroups and gyrocommutative gyrogroups proposed in UngarBeyonEinstein02; UngarAnalyticHyp05; UngarHyperbolicNDim. For greater mathematical detail and in-depth discussion, we refer the interested reader to these papers.
Definition G.1 (Gyrogroups (UngarHyperbolicNDim)).
@@
$a\oplus(b\oplus c)=(a\oplus b)\oplus\operatorname{gyr}[a,b]c\hskip 8.53581pt\text{(Left Gyroassociative Law)}.$
@@
Definition G.2 (Gyrocommutative Gyrogroups (UngarHyperbolicNDim)).
@@
$a\oplus b=\operatorname{gyr}[a,b](b\oplus a)\hskip 8.53581pt\text{(Gyrocommutative Law).}$
The following definition of gyrovector spaces is slightly different from Definition 3.2 in UngarHyperbolicNDim.
@@ G.2 AI Gyrovector Spaces
For $\mathbf{P},\mathbf{Q}\in\operatorname{Sym}_{n}^{+}$, the binary operation (NguyenECCV22) is given as
@@
The inverse operation (NguyenECCV22) is given by
@@ G.3 LE Gyrovector Spaces
For $\mathbf{P},\mathbf{Q}\in\operatorname{Sym}_{n}^{+}$, the binary operation (NguyenECCV22) is given as
@@
The inverse operation (NguyenECCV22) is given as
@@ G.4 LC Gyrovector Spaces
For $\mathbf{P},\mathbf{Q}\in\operatorname{Sym}_{n}^{+}$, the binary operation (NguyenECCV22) is given as
@@
The inverse operation (NguyenECCV22) is given by
@@ G.5 Grassmann Manifolds in The Projector Perspective
For $\mathbf{P},\mathbf{Q}\in\operatorname{Gr}_{n,p}$, the binary operation (NguyenNeurIPS22) is given as
@@
where $[.,.]$ denotes the matrix commutator.
The inverse operation (NguyenNeurIPS22) is defined as
@@ G.6 Grassmann Manifolds in The ONB Perspective
For $\mathbf{U},\mathbf{V}\in\widetilde{\operatorname{Gr}}_{n,p}$, the binary operation (NguyenGyroMatMans23) is defined as
@@
The inverse operation can be defined using the approach in NguyenGyroMatMans23 (see Section 2.3.1), i.e.,
$\widetilde{\ominus}_{gr}\mathbf{U}=\tau^{-1}\big(\ominus_{gr}(\mathbf{U}\mathbf{U}^{T})\big),$
@@ G.7 The SPD and Grassmann Inner Products
where $\langle.,.\rangle_{\mathbf{I}_{n,p}}$ denotes the inner product at $\mathbf{I}_{n,p}$ given by the canonical metric of $\operatorname{Gr}_{n,p}$.
@@ G.8 The Gyrocosine Function and Gyroangles in Structure Spaces
where $\|.\|^{psd,g}$ is the norm induced by the inner product in structure spaces. The gyroangle $\alpha$ is denoted by $\alpha=\angle\mathbf{Q}\mathbf{P}\mathbf{R}$.
@@ Appendix H Computation of Canonical Representations
We first recall the definition of the binary operation $\oplus_{g}$ in NguyenNeurIPS22.
Definition I.1 (The Binary Operation (NguyenNeurIPS22)).
@@
The first part of Proposition 3.4 can be easily verified using the definition of the SPD inner product (see Definition G.4) and that of Affine-Invariant metrics (PennecRiemMedicalImageAnalysis20) (see Chapter 3).
To prove the second part of Proposition 3.4, we will use the notion of SPD pseudo-gyrodistance (NguyenGyroMatMans23) in our interpretation of FC layers on SPD manifolds, i.e., the signed distance is replaced with the signed SPD pseudo-gyrodistance in the interpretation given in Section 3.2.1. First, we need the following result from NguyenGyroMatMans23.
Theorem J.1 (The SPD Pseudo-gyrodistance from an SPD Matrix to an SPD Hypergyroplane in an AI Gyrovector Space (NguyenGyroMatMans23)).
@@
$\begin{split}v_{(i,j)}(\mathbf{X})&=\frac{\langle\log(\mathbf{Y}),\frac{1}{\sqrt{2}}(\mathbf{e}_{i}\mathbf{e}^{T}_{j}+\mathbf{e}_{j}\mathbf{e}^{T}_{i})\rangle_{F}}{\|\frac{1}{\sqrt{2}}(\mathbf{e}_{i}\mathbf{e}^{T}_{j}+\mathbf{e}_{j}\mathbf{e}^{T}_{i}) \|_{F}}\\ &=\langle\log(\mathbf{Y}),\frac{1}{\sqrt{2}}(\mathbf{e}_{i}\mathbf{e}^{T}_{j}+\mathbf{e}_{j}\mathbf{e}^{T}_{i})\rangle_{F}\\ &=\frac{1}{\sqrt{2}}\big(\log(\mathbf{Y})_{(i,j)}+\log(\mathbf{Y})_{(j,i)}\big)\\ &=\sqrt{2}\log(\mathbf{Y})_{(i,j)}.\end{split}$
@@
$\begin{split}v_{(i,i)}(\mathbf{X})&=\frac{\langle\log(\mathbf{Y}),\mathbf{e}_{i}\mathbf{e}^{T}_{i}-\frac{1}{m}\big(1-\frac{1}{\sqrt{1+m\beta}}\big)\mathbf{I}_{m}\rangle_{F}}{\|\mathbf{e}_{i}\mathbf{e}^{T}_{i}-\frac{1}{m}\big(1-\frac{1}{\sqrt{1+m\beta}}\big)\mathbf{I}_{m} \|_{F}}\\ &=\langle\log(\mathbf{Y}),\mathbf{e}_{i}\mathbf{e}^{T}_{i}-\frac{1}{m}\big(1-\frac{1}{\sqrt{1+m\beta}}\big)\mathbf{I}_{m}\rangle_{F}.\end{split}$
@@
The first part of Proposition 3.6 can be easily verified using the definition of the SPD inner product (see Definition G.4) and that of Log-Cholesky metrics (Lin_2019).
To prove the second part of Proposition 3.6, we first recall the following result from NguyenGyroMatMans23.
Theorem L.1 (The SPD Gyrodistance from an SPD Matrix to an SPD Hypergyroplane in a LC Gyrovector Space (NguyenGyroMatMans23)).
@@
[…]
@@
where $\|$ is the concatenation operation similar to operation $\operatorname{concat}_{spd}(.)$.
@@
Now, the problem in (9) is to find the minimum angle between the vector $\sqrt{\lambda}\mathbf{A}_{1}\|\mathbf{A}_{2}]$ and the Euclidean hyperplane described by Eq. (11). The pseudo-gyrodistance from $\mathbf{X}$ to $\mathcal{H}_{\mathbf{W},\mathbf{P}}^{psd,g}$ thus can be obtained as
@@
We need the following result from NguyenGyroMatMans23.
@@
[…]
We adopt the notations in bendokat2020grassmann. The Riemannian metric $g^{O}_{\mathbf{Q}}(.,.)$ on $\operatorname{O}_{n}$ is the standard inner product given (Edelman98; bendokat2020grassmann) as
@@
Let $\mathbf{U}\in\widetilde{\operatorname{Gr}}_{n,p}$, $\mathbf{D}_{1},\mathbf{D}_{2}\in T_{\mathbf{U}}\widetilde{\operatorname{Gr}}_{n,p}$. The canonical metric $g^{\widetilde{\operatorname{Gr}}}_{\mathbf{U}}(\mathbf{D}_{1},\mathbf{D}_{2})$ on $\widetilde{\operatorname{Gr}}_{n,p}$ is the restriction of the Riemannian metric $g^{O}_{\mathbf{Q}}(.,.)$ to the horizontal space of $T_{\mathbf{Q}}\operatorname{O}_{n}$ (multiplied by 1/2) and is given (Edelman98; bendokat2020grassmann) by
@@
Let $\mathbf{P}\in\operatorname{Gr}_{n,p}$, $\Delta_{1},\Delta_{2}\in T_{\mathbf{P}}\operatorname{Gr}_{n,p}$. The canonical metric $g^{Gr}_{\mathbf{P}}(\Delta_{1},\Delta_{2})$ on $\operatorname{Gr}_{n,p}$ is the restriction of the Riemannian metric $g^{O}_{\mathbf{Q}}(.,.)$ to the horizontal space of $T_{\mathbf{Q}}\operatorname{O}_{n}$ (multiplied by 1/2) and is given (Edelman98; bendokat2020grassmann) by